OpenStudy (anonymous):

Can anyone help me solve this? 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = quantity n times quantity six n squared minus three n minus one all divided by two

OpenStudy (anonymous):

I already found n=1. i'm just having some trouble with n=k n=k+1

OpenStudy (anonymous):

i solved the left side i just don't know what to do with that fraction on the right

OpenStudy (mokeira):

ok

OpenStudy (anonymous):

@Mokeira do you think you can help me please?

OpenStudy (mokeira):

what is in between 7^2 + ... + (3n - 2)^2

OpenStudy (anonymous):

there's nothing in between those just those ellipses

OpenStudy (anonymous):

i set it up to n=1 and got true

OpenStudy (mokeira):

ok..let me try it

OpenStudy (anonymous):

okie doks :)

OpenStudy (mokeira):

does the quantity n multiply six n squared or everything?

OpenStudy (anonymous):

I'll set it up sorry about that

OpenStudy (anonymous):

(n(6n^2-3n-1))/2 that's the right side

OpenStudy (mokeira):

ok..great!

OpenStudy (anonymous):

@dg2

OpenStudy (anonymous):

but idk if i should add the fractions together

OpenStudy (anonymous):

@ApollosChariot

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

formula i think 1/6n(n+1)(2n+1)

OpenStudy (anonymous):

there's no formula...i got up to here:6k^3-3k^2-k/2 +(3k+1)^2

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

that's not how you do it tho but thanks for trying I appreciate it

OpenStudy (midhun.madhu1987):

This is related to Principle of mathematical induction. The various steps to prove it are: 1. Prove it for n=1. 2. Assume it is true for n=k. 3. Then we prove n=k+1 for every n=k. As an example, please find the attached file written by me to solve these kinds of questions... let me know if you need any help or clarifications...

OpenStudy (anonymous):

okay thanks I finally understand

OpenStudy (mokeira):

hey @midhun.madhu1987 when you are proving for P(1) you only use n, why dont you use 2+4+6?

OpenStudy (midhun.madhu1987):

Because if the value of n=1, then it contains only one term.. that is 2n.

OpenStudy (mokeira):

yeah...i have understood :)